On the first day of school, the percentage of boys in a particular class is 60%. During the school year, six girls move away, and are replaced in the class by six boys; this makes the class roster 75% boys. Find the number of boys and girls in the class on the first day of school.
Answer by checkley71(8403)

.6X=BOYS
.6X+6=.75X
.6X-.75X=-6
-.15X=-6
X=-6/-.15
X=40 CLASS SIZE ON THE FIRST DAY OF SCHOOL.
.6*40=24 BOYS ON THE FIRST DAY OF SCHOOL.
40-24=16 GIRLS ON THE FIRST DAY OF SCHOOL.
PROOF
.6*40+6=.75*40
24+6=30
30=30